OpenStudy (anonymous):

i tried this problem twice and both times got a different answer please help |3-5|-|-6|8-4|/|8-5|

OpenStudy (anonymous):

can you re-write it?

OpenStudy (anonymous):

how

OpenStudy (anonymous):

after -6 and between 8-4 i dont know what signs are there

OpenStudy (anonymous):

oh my bad oops |3-5|-|-6|-|8-4|/|8-5| is that better

OpenStudy (anonymous):

(|-2|-|-6|-|4|)/(|3|) = (2-6-4)/(3) = -8/3

OpenStudy (anonymous):

okay that is what i got the second time the 1st time i got -17/3 not sure how i got that but i used calculators both times and got two different answers thanks